You are an industry analyst that specializes in an industry where the market inverse demand is P = 250 - 6Q. The external marginal cost of producing the product is MCExternal = 8Q, and the internal cost is MCInternal = 14Q.

Instructions: Enter your responses rounded to the nearest two decimal places.

a. What is the socially efficient level of output?

units


b. Given these costs and market demand, how much output would a competitive industry produce?

units


c. Given these costs and market demand, how much output would a monopolist produce?

units


d. Which of the following are actions the government could take to induce firms in this industry to produce the socially efficient level of output.

Instructions:  In order to receive full credit, you must make a selection for each option. For correct answer(s), click the box once to place a check mark. For incorrect answer(s), click twice to empty the box.

  • Pollution taxesunanswered
  • Nonrival consumptionunanswered
  • Pollution permitsunanswered
Answer #1

a) Social MC = 8Q+14Q = 22Q

250-6Q = 22Q

250 = 28Q

Q = 250/28 = 8.93

b) Internal MC = MR=AR=P

14Q = 250-6Q

20Q = 250

Q = 250/20 = 12.5

c) MC = MR

MR = 250-12Q

14Q = 250-12Q

26Q = 250

Q = 250/26 = 9.62

d) Correct=

  • Pollution taxes
  • Pollution permits